TEAM OF 5 IS TO BE CHOSEN FROM 4 MEN AND 5 WOMEN TO WORK ON SPECIAL PROJECT.
1. In how many ways can the team be chosen?
2.In how many ways can the team be chosen to include just 3 women?
3.what is the probability that the team includes at least 3 women?
4.What is the probability that the team includes more men than women

Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

total number of ways is 9C5=126
with 3 women it would be 5C3*4C2=60 ways
at least 3 women is 60 (for 3) 5C4*4C1 or 20 (for 4) and 1 for 5, 81 ways for at least 3 women.
more men than women is at least 3 men.
that is 4C3*5C2=40 ways plus 4C4*5C1=5 ways
All this add to 126 and include all the possibilities.

Answer by Alan3354(69443)

TEAM OF 5 IS TO BE CHOSEN FROM 4 MEN AND 5 WOMEN TO WORK ON SPECIAL PROJECT.
1. In how many ways can the team be chosen?
If all positions are equal: --> 9*8*7*6*5
-------------------------
2.In how many ways can the team be chosen to include just 3 women?
5*4*3*4*3
---------------------------
3.what is the probability that the team includes at least 3 women?
4.What is the probability that the team includes more men than women